# A $2$-distance $(2\Delta+7)$-coloring of planar graphs

### v2

@@ 2.3. The case $\Delta=8$
(6.4). Let $m_{3}(v)=8$. By Lemma 2.34(b), $v$ has at most five 5(5)555(5)5 ( 5 )-neighbours. If $v$ has at most four such neighbours, then $\mu^{*}(v)\geq 4-8\times\frac{1}{3}-4\times\frac{2}{9}-4\times\frac{1}{9}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{2}{9}$ to each of its 5(5)555(5)5 ( 5 )-neighbours by R9(b), and at most $\frac{1}{9}$ to each of its $4(3)$- and $5(4)$-neighbours by R6(c) and R8. If $v$ has five 5(5)555(5)5 ( 5 )-neighbours, then $v$ has neither $4$-neighbour nor $5(4)$-neighbour by Lemma 2.34(a). Thus, $\mu^{*}(v)\geq 4-8\times\frac{1}{3}-5\times\frac{2}{9}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, and $\frac{2}{9}$ to each of its 5(5)555(5)5 ( 5 )-neighbours by R9(b).

### v3

@@ 2.3. The case $\Delta=8$
By Lemma 2.31(c), a $4(3^{+})$-vertex has no $4$- and $5(4^{+})$-neighbours. So, we deduce that $v$ has at most four $4(3^{+})$-neighbours. Indeed, if $v$ has four such neighbours, then $v$ is not adjacent to any $4(2^{-})$- or $5(4^{+})$-vertices by Lemma 2.31(c). Thus $\mu^{*}(v)\geq 4-7\times\frac{1}{3}-4\times\frac{1}{3}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{3}$ to each of its $4(3^{+})$-neighbours by R9(c). We may further assume that $v$ has at most three $4(3^{+})$-neighbours.
Suppose first that $v$ has no $4(3^{+})$-neighbour. Note that $v$ has at most six neighbours consisting of $4(2^{-})$- or $5(5)$-vertices by Lemma 2.40(g). Moreover, if $v$ has six such neighbours, then $v$ cannot be adjacent to any $5(4)$-vertex by Lemmas 2.31 and 2.37, and so $\mu^{*}(v)\geq 4-7\times\frac{1}{3}-6\times\frac{1}{4}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, at most $\frac{1}{4}$ to each of its $4(1)$-, $4(2)$- and $5(5)$-neighbours by R9(a), (b), (e). If $v$ has at most five neighbours consisting of $4(2^{-})$- or $5(5)$-vertices, then $\mu^{*}(v)\geq 4-7\times\frac{1}{3}-5\times\frac{1}{4}-3\times\frac{1}{9}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, at most $\frac{1}{4}$ to each of its $4(1)$-, $4(2)$- and $5(5)$-neighbours by R9(a),(b),(e), and $\frac{1}{9}$ to each of its $5(4)$-neighbours by R9(d).
Next we suppose that $v$ has exactly one $4(3^{+})$-neighbour. By Lemma 2.31(c), $v$ has at least one neighbour different from $4$- and $5(4^{+})$-vertices, which does not receive any charge from $v$. On the other hand, a $4(t)$-vertex with $1\leq t\leq 2$ has at most one neighbour consisting of $4$- or $5(5)$-vertices by Lemma 2.31(a),(b). It then follows from Lemma 2.31(c) and Lemma 2.37(b) that $v$ has at most four neighbours consisting of $4(1)$-, $4(2)$- or $5(5)$-vertices. Thus $\mu^{*}(v)\geq 4-7\times\frac{1}{3}-\frac{1}{3}-4\times\frac{1}{4}-2\times\frac{1}{9}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{3}$ to each of its $4(3^{+})$-neighbours by R9(c), at most $\frac{1}{4}$ to each of its $4(1)$-, $4(2)$- and $5(5)$-neighbours by R9(a),(b),(e), and $\frac{1}{9}$ to each of its $5(4)$-neighbours by R9(d).
Suppose that $v$ has exactly two $4(3^{+})$-neighbours. By Lemma 2.31(c), a $4(3^{+})$-vertex has neither $4$-neighbour nor $5(4^{+})$-neighbour. This means that $v$ has at least two neighbours different from $4$- and $5(4^{+})$-vertices, which do not receive any charge from $v$. Thus $\mu^{*}(v)\geq 4-7\times\frac{1}{3}-2\times\frac{1}{3}-4\times\frac{1}{4}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{3}$ to each of its $4(3^{+})$-neighbours by R9(c), and at most $\frac{1}{4}$ to each of its $4(1)$-, $4(2)$- and $5(4^{+})$-neighbours by R9(a),(b),(d),(e).
Finally suppose that $v$ has exactly three $4(3^{+})$-neighbours. It follows from Lemma 2.31(c) that $v$ has at least three neighbours different from $4$- and $5(4^{+})$-vertices, which do not receive any charge from $v$. Thus $\mu^{*}(v)\geq 4-7\times\frac{1}{3}-3\times\frac{1}{3}-2\times\frac{1}{4}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{3}$ to each of its $4(3^{+})$-neighbours by R9(c), at most $\frac{1}{4}$ to each of its $4(1)$-, $4(2)$- and $5(4^{+})$-neighbours by R9(a),(b),(d),(e).
(6.4). Let $m_{3}(v)=8$. By Lemma 2.37(b), $v$ has at most five $5(5)$-neighbours. If $v$ has at most four such neighbours, then $\mu^{*}(v)\geq 4-8\times\frac{1}{3}-4\times\frac{2}{9}-4\times\frac{1}{9}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{2}{9}$ to each of its $5(5)$-neighbours by R10(b), and at most $\frac{1}{9}$ to each of its $4(3)$- and $5(4)$-neighbours by R10(a). If $v$ has five $5(5)$-neighbours, then $v$ has neither $4$-neighbour nor $5(4)$-neighbour by Lemma 2.37(a). Thus, $\mu^{*}(v)\geq 4-8\times\frac{1}{3}-5\times\frac{2}{9}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, and $\frac{2}{9}$ to each of its $5(5)$-neighbours by R10(b).
@@ Acknowledgement
Declarations
Conflict of interest The author has no conflicts of interest to declare that are relevant to the content of this article.
Availability of data and material Data sharing not applicable to this article as no datasets were generated or analysed during the current study.